OpenStudy (anonymous):

multiply and simplify (2√7+3√5)(√7-2√5)

OpenStudy (anonymous):

ok multiply one by one by removing the bracket and using the distributive property

OpenStudy (anonymous):

\[({2}\sqrt{7} + {3}\sqrt{5})({\sqrt{7}}-2\sqrt{5})\] \[\sqrt{7}\times({2}\sqrt{7} + {3}\sqrt{5})-2\sqrt{5}\times({2}\sqrt{7} + {3}\sqrt{5})\]

OpenStudy (anonymous):

so i got -√35-52

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

14-4√35+3√35-30 am i right so far